# thanks: Corresponding author

### v1

@@ II Ising Model on random graph
$\{X_{i},Y_{i},Z_{i}\}$ are Pauli matrices acting on site $i$ and $g,h$ are the strengths of the transverse and longitudinal field, respectively. $A_{ij}$ is the adjacency matrix for the underlying graph where spins are placed on nodes. $A_{ij}=1(0)$ if there is an edge (no edge) between nodes $i$ and $j$. Here, edges between two nodes represent interaction between the spins placed on them. We take $J=1$ throughout our discussion. The adjacency matrices are chosen to represent random Erdős-Rényi graphs [44] $\mathcal{G}(L,M)$ where $L$ and $M$ denote the number of nodes and edges, respectively. $M$ can be obtained as, $M=\frac{1}{2}\sum_{i,j}A_{ij}$. The denominators $L,M$ in the Hamiltonian ensure that the bandwidth of the energy spectrum scales as $\approx O(JL^{0}M^{0})$. For $g=0$, there is a global $U(1)$ symmetry. For generic $g$, at $h=0$, the model has a global parity symmetry given by, $\mathcal{P}=\prod_{i=1}^{L}X_{i}$. All edges have equal strength ($J$) and graph-to-graph variation arises entirely from the adjacency matrix. For $L$ sites, the total number of possible edges is $\binom{L}{2}$. When $M$ edges are selected, there are $\binom{N_{\text{max}}}{M}$ possible random graphs where $N_{\text{max}}=\binom{L}{2}$. We sample these with equal probability and define the connectance $\tilde{M}=M/N_{\text{max}}$ to characterize the lattice graphs in subsequent discussion. $\tilde{M}\in[0,1]$ where $\tilde{M}=0,1$ corresponds to non-interacting and integrable LMG [37] limit respectively (the presence of small longitudinal field causes deviation in spectral characteristics from the LMG model in finite systems). Figure 1 illustrates a single realization of such a lattice. We take $J=1,g=1$ and $h$ to be a small positive value, thus working with a ferromagnetic Hamiltonian. The dynamical properties studied in this work characterize the bulk of the spectrum, which is expected to be independent of the ferromagnetic or anti-ferromagnetic nature of the ground state. Note that the QAA usually targets the edge of the spectrum (ground state).
The onset of chaos with $\tilde{M}$ in this model (with $h=0$) has been studied [32] through spectral analysis. The chaotic behavior sets in for intermediate $\tilde{M}$ and the spectrum shows localization for extreme $\tilde{M}$. For finite $L$, the system appears integrable at $M\to 0,1$. Through spectral analysis and perturbative arguments, it was shown that the distribution of the level-spacing in the unfolded spectrum converges to the Wigner-Dyson(WD) distribution at large $L$ for $\tilde{M}\lesssim 1$ and $\tilde{M}\gtrsim 0$ for intermediate $g(0.5J<g<4J)$. At very large $g$, the system is effectively non-interacting with the eigenstates given by simple product states; thus, the level-spacing ratio starts deviating from WD. At extreme $\tilde{M}$, the level spacing ratio resembles a Poisson distribution, signaling the breakdown of quantum chaos.
We choose $g=1$ throughout the discussion to retain the possibility of chaotic behavior in the system sizes studied ($6\leq L\leq 14$). We will take $h=0.1$ to break the parity symmetry unless mentioned otherwise. Having established the Hamiltonian, we first investigate its thermalization properties at the subsystem level using the projected ensemble and the partial spectral form factor. We then turn to the system’s dynamical response, analyzing the relaxation of both operators and states across the range of connectance, $\tilde{M}$.
III Projected ensemble and emergence of deep thermalization with $\tilde{M}$
@@
At the level of the density matrix, the ETH predicts that subsystems of unitarily evolving isolated systems relax to Gibbs ensembles determined by the initial state. This relaxation emerges from unitary evolution scrambling initially localized information across the entire system, with the subsystem’s complement serving as an effective bath whose degrees of freedom are traced out to obtain the reduced density matrix. However, modern quantum simulator platforms can precisely track all degrees of freedom of the system simultaneously [45, 46, 47], thus making the distinction between the system and bath artificial. Access to the state of the bath allows us to obtain conditional states in the subsystem conditioned on the measurement outcome of the bath. The ensemble of these conditional states is denoted as the Projected Ensemble, which can be considered as a natural unraveling of the reduced density matrix of the subsystem. Unlike the density matrix, which captures only the first moment of the state ensemble, the PE encodes the full distribution of quantum states in Hilbert space, providing access to higher moments and fluctuations of observables beyond their expectation values. This perspective naturally extends the question of quantum thermalization beyond local observables to whether the underlying distribution of quantum states also approaches a universal form. This phenomenon of relaxation of the quantum state distribution, dubbed “deep thermalization”, generalizes the notion of thermalization beyond the density matrix and expectation value of local observables.
To characterize deep thermalization, it is essential to identify the universal ensemble toward which the PE is expected to converge. In the absence of any conservation law, the Haar ensemble [48, 49] is the maximal entropic ensemble of pure quantum states in Hilbert space. For a system initialized in an effective infinite temperature state and evolved under a chaotic Hamiltonian, the subsystem’s reduced density matrix asymptotically approaches the identity, while the moments of the PE simultaneously approach those of the Haar ensemble [38, 39]. The proximity of the PE to the Haar ensemble serves as a measure of deep thermalization. This can be quantified in terms of the distance between the corresponding moments of the two ensembles. Given an ensemble of quantum states $\mathcal{E}=\{P(\psi),\psi\}$, their $k$-th moment is defined as $\rho_{\mathcal{E}}=\sum_{\psi}P(\psi)(|\psi\rangle\langle\psi|)^{\otimes k}$ where $P(\psi)$ is the probability of state $\psi$ in $\mathcal{E}$. The onset of deep thermalization has been explored in various settings including random circuits, generic Hamiltonian system and the systems where ETH breaks down [50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. In the presence of conservation laws or constraints, the asymptotic projected ensemble reached at $L,t\to\infty$ gets modified to the Scrooge ensemble, which is determined by the conserved charges in the initial state and measurement direction [56, 52].
Here, we briefly outline the protocol to construct PE. Consider a system of $L$ sites partitioned into $A$ and $B$ subsystems with $L=L_{A}+L_{B}$. The entire system is described by a pure state, $\psi_{AB}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ at all times. We repeat a set of projective measurements restricted to the subsystem $B$ on copies of $\psi_{AB}$ and collect the corresponding pure state in $A$ conditioned on the measurement outcome in $B$, $\psi_{A}(b_{i})$
$\psi_{A,b_{i}}=\sum_{b_{i}}\frac{1}{\sqrt{p(b_{i})}}\Pi_{b_{i}}\psi_{AB}=\psi_{A}(b_{i})\otimes|b_{i}\rangle,$ (2)
where $\Pi_{b_{i}}=\mathbb{I}\otimes|b_{i}\rangle\langle b_{i}|$ and $p(b_{i})=\langle\psi_{AB}|\Pi_{b_{i}}|\psi_{AB}\rangle$. $\psi_{A}(b_{i})$ is the conditional pure state obtained in subsystem $A$ provided that the measurement result in $B$ is $b_{i}$. The PE is constructed as, $\mathcal{E}_{\mathrm{PE}}=\{p(b_{i}),\psi_{A}(b_{i})\}$. We can characterize the PE through its $k$-th moment as, $\rho^{(k)}_{\mathrm{PE}}=\sum_{b_{i}}p(b_{i})(|\psi_{A}(b_{i})\rangle\langle\psi_{A}(b_{i})|)^{\otimes k}$. Any moment of observables can be obtained with access to $\mathcal{E}_{\mathrm{PE}}$ as $\langle\hat{O}^{\otimes k}\rangle_{\mathrm{PE}}=\mathrm{Tr}[\rho^{(k)}_{\mathrm{PE}}\hat{O}^{\otimes k}]$. The reduced density matrix of the subsystem (and the expectation values of local observables obtained from it) corresponds to $\rho^{(1)}$, which is described by the ETH for chaotic systems at equilibrium. To compare two state ensembles, we define the trace distance between the $k$-th moments of two ensembles $\mathcal{E}$ and $\mathcal{E^{\prime}}$ as,
@@
Overbar denotes quantities averaged over multiple graph realizations. Quantum information is maximally scrambled by chaotic evolution, and we consider that the initial state corresponds to an effective infinite temperature state. In that case, the projected ensemble should approach the Haar ensemble[38, 52], the uniform distribution over all pure states in Hilbert space i.e. achieving maximum entropy in the quantum state distribution. For a system $A$, the Haar distribution as a measure on normalized quantum states $\psi$ can be written as,
[…]
where $d_{A}$ is the dimension of the Hilbert space of subsystem $A$. We consider the proximity of the projected ensemble and Haar ensemble as a metric for chaos and thermalization in this section.
We take initial states in the direct product form as given below,
$|\psi\rangle=\frac{1}{\sqrt{2^{L}}}\otimes_{i=0}^{L}(|0\rangle+\imath|1\rangle)$ (5)
All projective measurements in $B$ are performed in a Pauli-$Z$ basis. A corresponding inverse temperature $\beta$ can be associated with the pure state given as, $E(\psi)=\langle\psi|H|\psi\rangle\propto\mathrm{Tr}(He^{-\beta H)}$. For a traceless Hamiltonian (like in Eq. 1), the mean of the energy spectrum is 0 i.e., zero-energy states correspond to the middle of the eigen spectrum with $\beta=\infty$. The initial state taken in Eq. 5 is such a zero-energy state which is time-evolved with the Hamiltonian in Eq. 1 with h=0.1,g=1h=0.1,\ g=1. Since there are no other conserved charges, the reduced density matrix of a subsystem corresponds to an infinite-temperature density matrix given as the identity. Note that the choice of other measurement bases does not qualitatively affect the convergence of PE in this setup due to the absence of conserved charges in the chaotic regime [56, 52].

### v2

@@ II Ising Model on random graphs
$\{X_{i},Y_{i},Z_{i}\}$ are Pauli matrices acting on site $i$ and $g,h$ are the strengths of the transverse and longitudinal field, respectively. $A_{ij}$ is the adjacency matrix of the underlying graph, where spins are the nodes. $A_{ij}=1(0)$ if there is a coupling (no coupling) between spins $i$ and $j$. We take $J=1$ throughout this work. The adjacency matrices are chosen to represent random ER graphs [37] $\mathcal{G}(L,M)$ where $L$ and $M$ denote the number of nodes and edges, respectively. $M$ can be obtained as $M=\frac{1}{2}\sum_{i,j}A_{ij}$. The denominators $L,M$ in Eq. 1 ensure that the bandwidth of the energy spectrum scales as $\approx O(JL^{0}M^{0})$. For $g=0$, there is a global $U(1)$ symmetry. For generic $g$, at $h=0$, the model has a global parity symmetry given by, $\mathcal{P}=\prod_{i=1}^{L}X_{i}$. All edges have equal strength ($J$) and graph-to-graph variation arises entirely from the adjacency matrix. For $L$ sites, the maximum possible number edges is $M_{\text{max}}=\binom{L}{2}$. We define the connectance $\tilde{M}=M/M_{\text{max}}$ to characterize the graphs. For each $\tilde{M}$ we sample graphs with $\tilde{M}M_{\rm max}$ edges with uniform probability.
Figure 1: (a) A single realization of Erdos-Renyi graph with $L=8$ nodes and connectance, $\tilde{M}=0.6$ which implies number of connected edges, $M=17$. Interacting pairs are represented by connected edges. (b) Eigen-spectrum of Hamiltonian in Eq. 1 for an instance at each $\tilde{M}$ with $L=8$. We have used darker shades to indicate degeneracy.
$\tilde{M}=0$ and $1$ correspond to the non-interacting and integrable LMG [61] limits, respectively (the presence of a small longitudinal field causes deviation in spectral characteristics from the LMG model in finite systems). Figure 1(a) shows one realization of such a lattice. We take $J=1$ and $g=1$, thus working with a ferromagnetic Hamiltonian. $h$ is taken to be a small positive value in subsequent sections to break the model’s parity symmetry. The dynamical properties studied in this work characterize the bulk of the spectrum, which is expected to be independent of the ferromagnetic or anti-ferromagnetic nature of the ground state, i.e., the sign of $J$.
The onset of chaos with $\tilde{M}$ in this model (with $h=0$) was studied in Ref. [42] using level spacing statistics, level velocity, and spectral form factor. The system exhibits chaotic behavior at intermediate connectivity, with spectral features matching Wigner-Dyson (WD) statistics, and shows deviations at extreme $\tilde{M}$ in finite-system studies. Figure 1(b) shows the eigenvalue spectrum of this Hamiltonian at $L=8,g=1.0$ and $h=0.1$ for a representative realization at each $\tilde{M}$. For intermediate $\tilde{M}$, the level-spacing distribution in the unfolded spectrum is close to the WD distribution for numerically accessible systems. It approaches the WD distribution with increasing $L$ for $\tilde{M}$ close to $1$ and 0 when $g$ takes a value in $0.5J<g<4J$. Thus, at large enough $L$, and intermediate values of $g$, the level spacing ratio is expected to be WD for all $\tilde{M}\neq 0,1$. At very large and small values of $g$, the system is nearly non-interacting, with eigenstates given by simple product states to a good approximation; thus, the level-spacing ratio deviates from WD.
We choose $g=1$ throughout to retain the possibility of chaotic behavior in the system sizes studied ($6\leq L\leq 15$). Next, we discuss the symmetries of this Hamiltonian at extremal $\tilde{M}$.
II.1 Symmetries and conservation laws near $\tilde{M}=0$ and 11
At small $\tilde{M}$ and finite system sizes, ER graphs consist of weakly connected clusters. The Hamiltonian then decomposes effectively into smaller blocks with support only within individual clusters, leading to a breakdown of ergodicity. Consequently, eigenstates in this regime are weakly entangled, with bipartite entanglement entropy across a real-space cut remaining well below the volume-law value expected for chaotic systems, a trend we confirm for $\tilde{M}<0.3$ (Fig. 14 in Appendix A). However, the percolation threshold for ER graphs is $\tilde{M}_{c}\sim 1/L$, so for any fixed small $\tilde{M}>0$, increasing the system size $L$ eventually drives the graph past this threshold, beyond which a giant component of size $\mathcal{O}(L)$ emerges [38]. Equivalently, at fixed $L$, chaotic behavior sets in once $\tilde{M}$ exceeds $\mathcal{O}(1/L)$. This structural crossover in the $(\tilde{M},L)$ plane underlies the system-size-dependent onset of chaos observed in our study at small $\tilde{M}$.
At $\tilde{M}=1$ and $h=0$, the Hamiltonian in Eq. 1 can be rewritten as
$H=\frac{1}{2M}S_{z}^{2}+\frac{1}{L}S_{x},$ (2)
up to an additive constant, where $S_{x}=\sum_{i}X_{i}$ and $S_{z}=\sum_{i}Z_{i}$. The Hamiltonian is symmetric under the permutation group $S_{L}$ (relabeling of spin indices) and commutes with the total spin squared $S^{2}\equiv S_{x}^{2}+S_{y}^{2}+S_{z}^{2}$. As a consequence, the vector space can therefore be decomposed into
$V=\sum_{S={S_{\rm min}}}^{L/2}V_{S}\otimes V_{P},$ (3)
where $S_{\rm min}=0(1/2)$ if $L$ is even (odd). $V_{S}$ forms a spin $S$ irrep of $SU(2)$ and $V_{P}$ is some irrep of $S_{L}$ representing the multiplicity of the $S$ representation. On this decomposition, $H$ acts as $\sum_{S}^{\oplus}H_{S}\otimes\mathbb{I}_{V_{P}}$. To explicitly see this, we can use the following three-spin/qubit example. The space decomposes into irreps of $SU(2)$ as $\frac{3}{2}\oplus\frac{1}{2}\oplus\frac{1}{2}$. In the following basis
$|\frac{3}{2},m-\frac{3}{2}\rangle\propto S_{+}^{m}|000\rangle;\,m=0,1,\dots 3$
$|\frac{1}{2},-\frac{1}{2},\alpha=0\rangle\propto(|01\rangle-|10\rangle)|0\rangle$
$|\frac{1}{2},+\frac{1}{2},\alpha=0\rangle\propto(|01\rangle-|10\rangle)|1\rangle$
$|\frac{1}{2},\frac{1}{2},\alpha=1\rangle\propto\sqrt{\frac{2}{3}}\,|110\rangle-\frac{1}{\sqrt{6}}(|101\rangle+|011\rangle)$
$|\frac{1}{2},-\frac{1}{2},\alpha=1\rangle\propto\frac{1}{\sqrt{6}}(|100\rangle+|010\rangle)-\sqrt{\frac{2}{3}}\,|001\rangle$
where the $\alpha$ index represents the multiplicity sector, $S_{+}$ is the collective spin raising operator, Hamiltonian has the form
$H=\left[\begin{array}[]{ccc}H_{\frac{3}{2}}&0&0\\ 0&H_{\frac{1}{2}}&0\\ 0&0&H_{\frac{1}{2}}\end{array}\right]\equiv H_{\frac{3}{2}}\oplus H_{\frac{1}{2}}\otimes\mathbb{I}_{2}$ (4)
In a similar way, the $L=4$ case decomposes as $H_{2}\oplus H_{0}\otimes\mathbb{I}_{2}\oplus H_{1}\otimes\mathbb{I}_{3}$.
The permutation symmetry and the commutation with $S^{2}$ lead to extensive Hilbert space fragmentation with maximal fragment size upper bound by $L+1$ (corresponding to the $S=\frac{L}{2}$ sector). Less important here, the permutation symmetry guarantees extensive degeneracy from the multiplicity space acting as a zero mode. The small fragments imply that the initial states undergo very restricted expansion in the Hilbert space, resulting in the absence of thermalization. As the bonds are removed when $\tilde{M}<1$, the symmetries are lost, resulting in weak coupling of the sectors. Eigenstates still have large weights in one block, resulting in an eigenstate ordering inherited from the nearby $\tilde{M}=1$ point.
When $\tilde{M}=1-\epsilon$ for small $\epsilon>0$, i.e., just below the maximally connected limit, the Hamiltonian can be viewed as a perturbation of the LMG model by a term
$V=\frac{2J}{L(L-1)\tilde{M}}\sum_{\langle ij\rangle_{\rm del}}Z_{i}Z_{j},$ (5)
where the sum runs over the $\epsilon L(L-1)/2$ deleted bonds. Before the perturbation, the LMG energy eigenvalues depend only on the $S^{2}$ eigenvalue, $E=E(s)$. The perturbation introduces off-diagonal blocks coupling different $s$ sectors. For $s$ away from the fully polarized limits, each $s$ sector is exponentially degenerate, with multiplicity
$m(s)=\frac{2s+1}{L/2+s+1}\binom{L}{L/2-s}$. (6)
Treating the degenerate eigenstates as random vectors in the computational basis with typical coefficient magnitude $\sim 2^{-L/2}$, the off-diagonal blocks coupling sectors $s$ and $s^{\prime}$ have dimensions $m(s)\times m(s^{\prime})$ (exponentially large in $L$) and matrix elements of order $\langle s|V|s^{\prime}\rangle\sim\mathcal{O}(\sqrt{\epsilon}/2^{L/2})$. Although individual matrix elements are exponentially small, the coupling between exponentially large degenerate blocks leads to a lifting of the block-diagonal structure and mixing of eigenstates at large $L$. While this constitutes only a plausibility argument for the emergence of chaos at any finite $\epsilon$, our numerical results are consistent with this picture.
While an atypical initial state, such as a direct product state, will have overlap with an extensive number of energy eigenstates, taking it to an infinite temperature diagonal ensemble at equilibration, the eigenstate ordering results in overlaps with only a restricted set of eigenstates, resulting in the diagonal ensemble deviating from the infinite temperature ensemble. We explicitly show this distinction through the inverse participation ratio (IPR) of such product state in the basis of eigenstates (Fig. 15 in Appendix A).
In the rest of this section, we demonstrate that chaotic dynamics in such a Hamiltonian can affect QAOA performance when this Hamiltonian is employed as an additional term in QAOA protocol.
II.2 QAOA with random Ising driver
Here, we empirically demonstrate, in the specific setting of QAOA, that chaos in the instantaneous Hamiltonian can be relevant to the circuit performance. We consider the problem of minimizing a classical Ising cost Hamiltonian of the form
$H_{\rm c}$ $=\sum_{i,j\in C}W_{ij}Z_{i}Z_{j},$ (7)
defined on a random graph $C$ with $L$ nodes. Several difficult classical optimization problems [7], such as MAXCUT and SAT, can be cast as ground-state search of such an Ising Hamiltonian [65]. For the numerical demonstrations in this section, we sample the graph $C$ from a set of ER graphs and sample the edge weights $W_{ij}$ uniformly from $\{+1,-1\}$. The standard QAOA ansatz alternates between the cost unitary $e^{\imath\gamma_{k}H_{\rm c}}$ and a mixer unitary $e^{\imath\beta_{k}H_{\rm x}}$ where $\gamma_{k},\beta_{k}$ are variational parameters associated with the $k$-th layer which are to be optimized. The entire circuit consists of $p$ such layers. Here, we modify this QAOA protocol by adding an additional driver unitary $e^{\imath\alpha_{k}H_{\rm d}}$ in each layer as illustrated in Fig. 2.
Figure 2: An outline of QAOA protocol. The three-term variational circuit is parameterized by $\{\alpha_{k},\beta_{k},\gamma_{k}\}$. The entire circuit consists of $p$ such layers. The input state of the ansatz is taken as $|+\rangle^{\otimes L}$. The energy of the cost function is optimized with a classical optimizer and the parameters are iteratively updated using gradient descent.
$H_{\rm d}$ is a mixed-field Ising model on an ER graph with Pauli-$YY$ interaction given as,
$H_{\rm d}^{YY}$ $=-\frac{JL}{M}\sum_{i,j}A_{ij}Y_{i}Y_{j}-g\sum_{i}X_{i}.$ (8)
Its spectral properties are identical to those of the annealing Hamiltonian (Eq. 1 with $h=0$ for simplicity), and its chaotic character is therefore tunable through $\tilde{M}$. Note that the adjacency matrices $C$ (of $H_{\rm c}$) and $A$ (of $H_{\rm d}$) are chosen independently. We take the initial state for the ansatz as $|+\rangle^{\otimes L}$ and set $J=g=1$. The values of $\alpha,\beta,\gamma$ are optimized such that $\langle H_{\rm c}\rangle$ is minimized. These unitaries need to be implemented on quantum devices using a Trotterized approach [36, 27].
To understand the state evolution in such a circuit, we examine the trajectory of the energy expectation value of the instantaneous state at each step of a 4-layer QAOA ansatz (Fig. 3). The background horizontal lines show the instantaneous eigenspectrum, colored by the overlap with the optimized state $|\psi_{\rm opt}\rangle$ ; the darkest lines, concentrated at the low-energy end of $H_{c}$, indicate that $|\psi_{\rm opt}\rangle$ is strongly supported on low-energy computational basis states. For optimized parameters (black line), the trajectory navigates through intermediate states with finite energy density, with significant overlap with mid-spectrum eigenstates of $H_{d}$ at intermediate steps. This is qualitatively distinct from adiabatic evolution, in which the state remains confined to the low-energy manifold at all times. Nevertheless, the optimized circuit successfully drives the final state to near-ground-state energy, as indicated by the diamond marker ($\langle\psi_{\rm opt}|H_{c}|\psi_{\rm opt}\rangle\approx-1$). For randomly chosen parameters (red dashed line), the energy expectation value shows no systematic drift toward low energy, fluctuating throughout the bulk of the instantaneous spectrum.
Figure 3: Trajectory of energy expectation value of instantaneous state in QAOA dynamics through the eigen-spectrum of instantaneous QAOA Hamiltonian. The thickness of dashes is proportional to the degeneracy of the eigenstates. The ansatz is a 44-layer quantum circuit. The label at the bottom of each stack shows the instantaneous Hamiltonian and the corresponding layer (in superscript). The marker in the right-most stack marks the final energy expectation value of the optimized state, i.e., $\langle\psi_{\rm{opt}}|H_{\rm c}|\psi_{\rm{opt}}\rangle$. The color indicates the overlap between the optimized state and the instantaneous eigenspectrum. $L=8$. The problem graph has a connectivity $\tilde{M}_{\rm c}=0.4$.
Figure 4: Approximation ratio of the optimized state for a QAOA circuit with $L=8,p=2$ (a,b) and $L=8,p=3$ (c,d). Keeping the problem $H_{\rm c}$ fixed, for each $\tilde{M}$, we show the QAOA optimized approximation ratio for $1000$ random choices of $H_{\rm d}$ at that $\tilde{M}$. The two columns describe data for two independently sampled problems $H_{\rm c}$ with $\tilde{M}_{\rm c}=0.4$.
We now assess the performance of this three-term QAOA protocol as a function of $\tilde{M}$. We show the results for two randomly picked instances of the classical Ising Hamiltonian $H_{\rm c}$ (Eq. 7) on a random ER graph of $L=8$ sites with $\tilde{M}_{\rm c}=0.4$. Keeping $H_{\rm c}$ fixed, we sample $1000$ different realizations of $H_{\rm d}$ (Eq. 8) and optimize a circuit of $p=2,3$ layers. In each case, $600$ ADAM [58] gradient-descent steps were performed to minimize the cost function $E_{\rm{QAOA}}=\langle\psi|H_{\rm c}|\psi\rangle$ (energy expectation value of the final state with respect to the cost Hamiltonian). $E_{\rm{QAOA}}$ is evaluated numerically exactly and the gradient is obtained using finite difference method. To gauge the distribution of performance of the optimized QAOA, we compute the approximation ratio for each of the $1000$ instances as defined below,
$\mathcal{R}=\frac{E_{\rm{max}}-E_{\rm{QAOA}}}{E_{\rm{max}}-E_{\rm{GS}}}\leq 1$ (9)
Higher $\mathcal{R}$ implies better optimization performance. Here, $E_{\rm{GS}}$ and $E_{\rm{max}}$ are the ground state and maximum eigenvalue of $H_{\rm c}$ respectively. Figure 4 shows the distribution of $\mathcal{R}$ values for the two cost Hamiltonians. Two features are particularly notable. First, $\mathcal{R}$ is higher at $\tilde{M}>0$ compared to the standard transverse-field mixer ($\tilde{M}=0$) for the majority of driver instances, confirming that the additional Ising mixer improves QAOA performance. Further, the approximation ratio is almost always larger at intermediate $\tilde{M}$, where $H_{\rm{d}}$ exhibits chaotic spectral statistics. Note that the spread in $\mathcal{R}$ at $\tilde{M}=1$ is due to the variation in initial parameter value across multiple gradient-descent optimizations. This correlation suggests possible influence of the chaotic dynamics of $H_{\rm{d}}$ on QAOA performance, similar to what was observed for QAA [90] (see Appendix B for additional numerical examples for system size up to $L=10$).
The improvement in QAOA performance observed with $H_{\rm d}$ containing $YY$ term can potentially be attributed to two distinct factors: the chaotic nature of $H_{\rm d}$, and the presence of terms that scatter between computational basis states with Hamming distance 2, enabling broader exploration of the solution space. To decouple these two contributions, we consider in Appendix B an alternative driver Hamiltonian with Pauli-$ZZ$ coupling,
$H_{\rm d}^{ZZ}$ $=-\frac{JL}{M}\sum_{i,j}A_{ij}Z_{i}Z_{j}-g\sum_{i}X_{i},$ (10)
in which the interaction term is diagonal in the computational basis but have identical spectral properties. We qualitatively obtain the same results, suggesting that the chaotic character of the entire $H_{d}$, rather than the nature of its individual terms, leads to improved performance of the modified QAOA.
Having demonstrated a potential connection between circuit performance and chaos in the dynamics, in the subsequent sections, we present a thorough investigation of the dynamical properties of $H_{\rm{d}}^{ZZ}$ utilizing multiple complementary probes.
III Projected ensemble and emergence of deep thermalization
@@
ETH predicts that the reduced density matrix of any small subsystem of an isolated quantum system evolving under a generic Hamiltonian relaxes to a Gibbs state determined solely by the initial energy density. This relaxation follows from the delocalization of initial local information across the many-body system, with scrambling as the underlying dynamical mechanism. The subsystem’s complement serves as an effective bath, whose degrees of freedom are traced out to yield the reduced density matrix. Under chaotic dynamics satisfying ETH, the resulting reduced state relaxes to a Gibbs ensemble, and all memory of the initial state is effectively erased from local observables. This framework is sufficient to predict expectation values of local subsystem observables. However, modern quantum simulator platforms can track all degrees of freedom simultaneously [5, 19, 74], rendering the system-bath distinction artificial. Access to the bath state allows one to obtain pure states of the subsystem conditioned on the measurement outcome of the bath. The ensemble of such conditional pure states is the Projected Ensemble (PE), and constitutes a physically natural unraveling of the reduced density matrix. While a density matrix admits infinitely many unravelings, the restriction to projective measurements selects a physically motivated one. Unlike the density matrix, which encodes only the first moment of the state ensemble, the PE encodes the full distribution of quantum states in the Hilbert space, providing access to all higher moments and thereby to fluctuations beyond mean observable values. This motivates a distinction between thermalization at the level of local expectation values and the more fundamental notion of deep thermalization, which concerns convergence of the underlying quantum state distribution in the projected ensemble to the Haar measure [66, 25, 70, 23, 49, 12, 69, 30, 104, 62].
Consider a system of $L$ sites partitioned into two subsystems, $A$ and $B$, with $L=L_{A}+L_{B}$. Under unitary evolution, the entire system is described by a pure state, $|\psi_{AB}\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ at all times. We consider initial states of the form
$|\psi_{AB}(t=0)\rangle=\frac{1}{\sqrt{2^{L}}}\bigotimes_{i=0}^{L}(|0\rangle+\imath|1\rangle),$ (11)
evolved under the Hamiltonian in Eq. 1 (with $J=1,g=1$ and $h=0.1$). The state has 0 energy expectation for this Hamiltonian. In the absence of any conserved charges (other than energy), the subsystem properties are expected to relax to those of a Gibbs ensemble $e^{-\frac{1}{T}H}$ where $T$ is the effective temperature of the initial state. As the Hamiltonian is traceless, $T=\infty$, and the subsystem properties of the time-evolved states (with Eq. 11 as initial state) relax to that of the identity density matrix.
We perform simultaneous projective measurements of Pauli-$Z$ on all qubits in subsystem $B$ on copies of $|\psi_{AB}\rangle$ and collect the corresponding pure state $|\psi_{A}(b_{i})\rangle$ in $A$ conditioned on the measurement outcome $b_{i}$,
$|\psi_{A,b_{i}}\rangle=\frac{1}{\sqrt{p(b_{i})}}\Pi_{b_{i}}|\psi_{AB}\rangle=|\psi_{A}(b_{i})\rangle\otimes|b_{i}\rangle,$ (12)
where $\Pi_{b_{i}}=\mathbb{I}\otimes|b_{i}\rangle\langle b_{i}|$ and $p(b_{i})=\langle\psi_{AB}|\Pi_{b_{i}}|\psi_{AB}\rangle$. The state $|\psi_{A,b_{i}}\rangle$ arises with probability $p(b_{i})$ and the ensemble of these states, described by $\mathcal{E}_{\mathrm{PE}}=\{p(b_{i}),|\psi_{A,b_{i}}\rangle\}$, is the PE. When seen as an ensemble of the states in $A$, its distribution over the Hilbert space of $A$ can be characterized by its $k$-th moment as, $\rho^{(k)}_{\mathrm{PE}}=\sum_{b_{i}}p(b_{i})(|\psi_{A}(b_{i})\rangle\langle\psi_{A}(b_{i})|)^{\otimes k}$. Given the PE, these can be computed using the $b_{i}$ labels on the ensemble states. The $k$th moment of any observable’s expectation value can be obtained as $\langle\hat{O}^{\otimes k}\rangle_{\mathrm{PE}}=\mathrm{Tr}[\rho^{(k)}_{\mathrm{PE}}\hat{O}^{\otimes k}]$.
To characterize deep thermalization, one must identify the universal ensemble toward which the PE converges in the limit of large system sizes $L$ and long times. Under unitary evolution with $H$ from the initial state in Eq. 11, if $H$ is chaotic, the PE asymptotically (in $L,t$) approaches the Haar ensemble [31, 28, 44, 109]. The choice of the measurement basis does not affect the convergence of the PE in this case [69, 70]. In the presence of conservation laws or constraints, the asymptotic projected ensemble reached at $L,t\to\infty$ deviates from the Haar ensemble and is determined by the charges of the initial state [69, 25, 70] as well as any correlation between measurement outcomes and the conserved charges. The degree of deep thermalization is thus quantified by the deviation of the PE moments from those of the Haar ensemble.
To compare the Haar ensemble and the PE, we use the trace distance between corresponding moments. Given two ensembles, we define the trace distance between their $k$-th moments as,
@@
We will denote quantities averaged over multiple graph realizations with overbar. The results presented here are averages over $500$ graph realizations. Figure 5 (a-d) show the evolution of trace distances $\Delta^{(1)}(\mathcal{E},\mathcal{E}_{\rm{Haar}})$ and $\Delta^{(2)}(\mathcal{E},\mathcal{E}_{\rm{Haar}})$ with time for different $\tilde{M}$. The distances decay with time in all cases, ultimately saturating to a system-size-dependent asymptotic value, $\overline{\Delta^{(k)}}(t\to\infty)$. Figure 6 shows the dependence of the asymptotic distance on the system size. Both the initial decay with time and the $L$ dependence of the asymptotic distance show clear dependence on $\tilde{M}$.